How do you subtract large mixed numbers like, 43 5/9 - 11 23/24 ?
Answer by CharlesG2(834)

How do you subtract large mixed numbers like, 43 5/9 - 11 23/24 ?

convert to fractions with common denominators

43 5/9 = 43 * 9/9 + 5/9 = 387/9 + 5/9 = 392/9
9 is not the same denominator as 24
3 * 24 = 72, 72 = 8 * 9
392/9 * 8/8 = 3136/72
11 23/24 = 11 * 24/24 + 23/24 = 264/24 + 23/24 = 287/24
287/24 * 3/3 = 861/72
3136/72 - 861/72 = 2275/72
72 is 8 * 9 and is only divisible by 2's and 3's
2275 is not divisible by 2 or 3
so 2275/72 can not be simplified, but it can be turned into a mixed number
31 * 72 = 2232
2275 - 2232 = 43
2275/72 = 31 43/72